# Learning Hierarchical Interactive Multi-Object Search for Mobile Manipulation

### v1

@@ Abstract
Existing object-search approaches enable robots to search through free pathways, however, robots operating in unstructured human-centered environments frequently also have to manipulate the environment to their needs. In this work, we introduce a novel interactive multi-object search task in which a robot has to open doors to navigate rooms and search inside cabinets and drawers to find target objects. These new challenges require combining manipulation and navigation skills in unexplored environments. We present HIMOS, a hierarchical reinforcement learning approach that learns to compose exploration, navigation, and manipulation skills. To achieve this, we design an abstract high-level action space around a semantic map memory and leverage the explored environment as instance navigation points. We perform extensive experiments in simulation and the real-world that demonstrate that HIMOS effectively transfers to new environments in a zero-shot manner. It shows robustness to unseen subpolicies, failures in their execution, and different robot kinematics. These capabilities open the door to a wide range of downstream tasks across embodied AI and real-world use cases.
@@ I Introduction
Autonomous navigation and exploration in unstructured indoor environments require a large variety of skills and capabilities. Pathways may be blocked and objects of interest may be stored away. Thus far, existing multi-object search tasks and methods have focused on environments that can be freely navigated with openly visible target objects [1, 2, 3]. We introduce a novel Interactive Multi-Object Search task in which target objects may be located inside articulated objects such as drawers and closed doors have to be opened to explore the environment. As a result, only navigation is insufficient to accomplish the task and the robotic agent has to physically interact with the environment to manipulate it to its needs.
@@
In this work, we propose Hierarchical Interactive Multi-Object Search (HIMOS), a hierarchical reinforcement learning approach to learn both exploration and manipulation skills and to reason at a high level about the required steps. We combine learned motions for local exploration in continuous action spaces [3] and frontier exploration for long-horizon exploration [4] together with mobile manipulation skills for object interactions [5]. We use semantic maps as the central memory component, which have shown to be an expressive and sample-efficient representation for these tasks [3] and design a high-level action space that exploits the acquired knowledge about the environment. By leveraging explored object instance locations as navigation waypoints, our approach efficiently learns these complex tasks from little data and consistently achieves success rates above 90% even as the number of target objects increases. By equipping all the low-level skills with mobility, we remove the ”hand-off” problem in which subpolicies have to terminate in the initial set of the following skill [6, 7]. Lastly, we transfer the trained agent to the real world and demonstrate that it successfully accomplishes these tasks in a real office environment. In particular, we replace the subpolicies from simulation with unseen real-world variations and find that the policy is able to generalize to these unseen subpolicies and is robust to failures in their execution, making it highly modular and flexible for transfer. Finally, we present ablation studies to evaluate the impact of the main design decisions.
@@
We present a hierarchical reinforcement learning approach that combines exploration and manipulation skills based on semantic knowledge and instance navigation points to efficiently solve these long-horizon tasks.
@@
We demonstrate the capabilities of this approach in both simulated and real-world experiments and show that it achieves zero-shot transfer to the real world, unseen environments, unseen subpolicies, and is robust to unseen failures.
@@ II Related Work
Object search tasks: Exploration and the ability to find items of interest is a key requirement for a wide range of downstream tasks. Previous work has proposed methods to maximize coverage of explored space [8] and to find specific objects of interest based on vision [9], auditory signals [10, 11] or target object categories [12, 13]. In ordered multi-object search tasks, the agent has to find items in a fixed order in game environments [14] or realistic 3D apartments [1]. In unordered multi-object search, the agent simply has to find the target objects as fast as possible, irrespective of the order [2, 3]. We focus on this unordered task. As we aim to demonstrate our system on a real robot, we follow Schmalstieg et al. [3] and use the full continuous action space. This is in contrast to most previous work which focuses on a simplified granular discrete action space. Existing search tasks assume that the desired goals can be freely reached by the agent. The interactive navigation task [15] relaxes this assumption by placing objects that the robot has to push away to reach the target. In contrast, we introduce an interactive search task for mobile robots equipped with a manipulator, that requires interaction with articulated objects to clear the path or reveal concealed objects. This requires integrating navigation and manipulation. Lastly, in contrast to most previous work, we demonstrate that our approach successfully transfers to the real world.
@@
Exploration requires both understanding and memorizing the seen environment and decision-making to explore the remaining space. Previous work has introduced both implicit and explicit memory mechanisms. Implicit memory agents either learn a direct end-to-end mapping from RGB-D images to actions or store embeddings of previous observations and retrieve them with an attention mechanism [2]. Other methods build explicit maps of the environment by projecting the RGB-D inputs into a global map. Commonly, this map is also annotated with semantic labels [14, 3, 1]. Further, combining short- and long-term exploration by learning an auxiliary prediction of the direction to the next closest object has proven to result in a strong performance in continuous action spaces [3]. We use this approach for low-level exploration. Frontier exploration [4] samples points on the frontier of the explored space and then navigate to these points. Instead of sampling, Ramakrishnan et al. [16] predict a potential function towards a target object. SGoLAM [17] combines mapping and a goal detection module. If no goal object is detected, it explores with frontier exploration and navigates directly to the goal otherwise. This results in strong results without any learning component. We include a similar exploration strategy as option in our hierarchical approach.
@@
Articulated object manipulation such as opening doors and drawers requires control of both the base and arm of the robot [18]. Existing approaches often separate both aspects and execute sequential navigation and manipulation. In our evaluation in simulation, we follow this approach and use BiRRT [19] to generate a motion plan for the robot arm. Recent work trains an agent to control the base of the robot via reinforcement learning to follow given end-effector motions [20, 5]. We use this method in our real world evaluation as it has been shown to generalize across different robots, tasks, and environments.
@@
Hierarchical methods introduce layers of abstraction by decomposing the decision-making into higher and lower-level policies. This shortens the time horizon of the Markov decision process (MDP) for the higher-level policies and enables the agent to combine different modules or skills at lower levels. At the same time, joint training of low- and high-level policies is often unstable and hard to optimize [21]. We focus on combining pretrained subpolicies. Joint finetuning of these policies offers a path to further performance improvements in the future. Pretraining is a common approach to increase the stability of the policy in hierarchical reinforcement learning [22]. While naive skill-chaining of arbitrary skills often results in ”hand-off” failures in which the subsequent skill cannot start from the current state [7], we resolve this issue by adding mobility to all the low-level skills in the real world execution, without the need for region-rewards [6]. A common navigation abstraction is to learn to set waypoints [23, 24], however, these often end up as very near points to the agent. In contrast, we propose instance navigation that provides a prior on important locations and action granularity. ASC [25] learns to combine combine navigation and pick skills for given receptacle locations. In contrast, we learn to search in unexplored environments with objects hidden in articulated objects. Alternatively, behavior trees are often used to decompose tasks hierarchically into a tree structure. These trees can either be fully constructed manually or be used in conjunction with a planner [26]. In contrast to these approaches, our proposed HIMOS learns high-level decision-making with a two-layer hierarchy.
@@ IV Hierarchical Interactive Multi-Object Search
The challenges introduced by this task require the robot to master an increasing number of different behaviors and skills. To stem this increase in complexity, we propose a hierarchical reinforcement learning approach and design efficient abstractions over states and actions. The model consists of three main components: a mapping module, a set of subpolicies for exploration, navigation, and interaction, and a high-level policy module to select the next low-level action to execute. An overview of the approach is depicted in Figure 2.
Assumptions: The focus of this work is on learning long-horizon decision-making, exploration, and search. We abstract from real-world perception and assume to have access to the following: (i) a semantic camera that produces accurate semantic object labels, (ii) accurate depth and localization, (iii) an object detection module to perceive object poses in the environment and to detect whether an object interaction was successful. In training and simulation, the object detection module uses ground truth poses from the simulator while in real-world experiments, we rely on AR markers placed on the respective objects. Interaction failure is detected based on the change in pose of these markers after the interaction.
@@ IV-A Mapping Module
A semantic map of the environment serves as a central memory component across all policies and modules. To construct this map, we build upon our previously introduced mapping module [3]. In the simulation, the robot receives the semantic labels from the simulator’s semantic camera, then uses the pixels depth and semantic camera observations to project the points into a local top-down map. In the real world, we simulate access to semantic labels as follows: we pre-build a map of the environment using Hector SLAM [27] and annotate it with labels for target objects, cabinets, and doors. All the other occupied space is mapped to the wall category. At test time, we then use the robot’s depth camera to build the same local map as in the simulation and overlay it with this pre-annotated semantic map. Localization in the real world is done on the same pre-recorded map. With the local map and its current localization, the agent then updates an internal global map which is further annotated with the agent’s trace and encoded into an RGB image.
In each step, the agent then extracts an egocentric map from the global map and passes two representations of this map to the encoder: a coarse map of dimension at a resolution of and a fine-grained map of dimension at a resolution of . I.e., they cover and , respectively. After the agent has segmented and approached a target object, it updates the object’s annotation to a fixed color coding to mark the corresponding object as “found”. We extend this map to a partial panoptic segmentation [28] map by labeling task-specific objects such as doors and cabinets with an instance-specific color, which is randomly assigned whenever a new instance is detected, i.e. an instance’s color changes from episode to episode, but remains consistent within an episode.
@@ IV-B Subpolicy Behaviors
The respective navigation goals are set to a pose slightly in front, or for the goal objects directly to the detected pose of the corresponding object. This navigation goal is then fed to an -planner which produces a trajectory at a resolution of . For training speed, we do not execute the full path in simulation, but set the robot’s base pose to the generated waypoints and only collect observations from these points. In the real world, we use the ROS navigation stack to move the robot to the goal. The policy may fail in some situations, for example due to collision with obstacles or narrow doorways. In this case, the agent returns to the last feasible waypoint, and given the updated semantic map , the high-level policy has to make a new decision.
@@ IV-C High-level Decision Making
Efficient high-level decision making requires the right level of abstractions of states and actions. We hypothesize that object- and instance-level decision making is such an efficient level of abstraction for embodied search tasks. We design a high-level policy around this idea. In particular: (i) We propose an instance navigation subpolicy that leverages the agent’s accumulated knowledge about the environment. It provides a prior on important places and on the granularity of navigation points, making it data-efficient and well-optimizable. (ii) As objects are discrete instances, the resulting full action space remains discrete, avoiding the complexities of mixed action spaces. At the same time, all the subpolicies still act directly in continuous action spaces, allowing for direct transfer to real robotic systems. (iii) We abstract from reasoning about exact robot placements in the real world by shifting the responsibility of mobility into the subpolicies. This ensures that the subpolicies can start from a large set of initial positions, resolving the ”hand-off” problem from naive skill-chaining [7] and strongly simplifies the learning process for the high-level policy. Furthermore, it enables us to change out the subpolicies to unseen subpolicies in the real world. (iv) We incorporate subpolicy failures into the training process, enabling the high-level policy to learn a retrial behavior if execution fails.
@@
The high-level policy acts in a Semi-Markov Decision Process (SMDP) in which the actions model temporarily extended behaviors and act at irregular intervals [30]. The high-level action space consists of (i) invoking the local exploration policy [3], (ii) invoking global frontier exploration [4], (iii) instance navigation with subsequent object interaction (if available), where instance IDs are selected through a one-hot vector mapping to fixed colors. The current task instantiation assumes a maximum of ten instances per episode, resulting in an overall 12-dimensional discrete action space.
@@
Adaptive Discounting: The high-level policy acts at irregular intervals, as the duration of the subpolicies varies largely. We correct for this time bias and accurately reflect the long-term consequences of actions with adaptive discounting [31].
@@
Invalid Action Masking: The availability of high-level actions varies with the state of the environment, e.g. navigating to an object instance depends on the instance being mapped, opening a cabinet is only possible if it is mapped and closed. We infer a valid actions vector from the agent’s observations and incorporate it into the training process by masking out invalid actions as well as including it in the observation space of the agent. As a result, the agent can learn more effectively, speeding up the training process [29]. We implement the masking by replacing the logits of the invalid actions with a large negative number.
@@
Architecture: The high-level policy observes the coarse semantic map which it encodes with a ResNet-18 [32]. Then it concatenates the features with the target object vector , the binary valid actions vector (see invalid action masking) and a robot state vector consisting of linear and angular base velocities, sum of collisions over the last ten steps, a current collision flag, and a normalized history over the last 16 high-level actions taken. The high-level policy is trained with Proximal Policy Optimization (PPO) [33], based on an open-source implementation [34]. We report hyperparameters and architecture details in the supplementary material.

### v3

# SReferences Learning Hierarchical Interactive Multi-Object Search for Mobile Manipulation

@@ Abstract
Existing object-search approaches enable robots to search through free pathways, however, robots operating in unstructured human-centered environments frequently also have to manipulate the environment to their needs. In this work, we introduce a novel interactive multi-object search task in which a robot has to open doors to navigate rooms and search inside cabinets and drawers to find target objects. These new challenges require combining manipulation and navigation skills in unexplored environments. We present HIMOS, a hierarchical reinforcement learning approach that learns to compose exploration, navigation, and manipulation skills. To achieve this, we design an abstract high-level action space around a semantic map memory and leverage the explored environment as instance navigation points. We perform extensive experiments in simulation and the real world that demonstrate that, with accurate perception, the decision making of HIMOS effectively transfers to new environments in a zero-shot manner. It shows robustness to unseen subpolicies, failures in their execution, and different robot kinematics. These capabilities open the door to a wide range of downstream tasks across embodied AI and real-world use cases.
@@ I Introduction
Autonomous navigation and exploration in unstructured indoor environments require a large variety of skills and capabilities. Pathways may be blocked and objects of interest may be stored away. Thus far, existing multi-object search tasks and methods have focused on environments that can be freely navigated with openly visible target objects [1, 2, 3]. However, as we move to human-centered environments such as households, neither of these assumptions hold. We introduce a novel Interactive Multi-Object Search task in which target objects may be located inside articulated objects such as drawers and closed doors have to be opened to explore the environment. As a result, only navigation is insufficient to accomplish the task and the robotic agent has to physically interact with the environment to manipulate it to its needs.
@@
In this work, we propose Hierarchical Interactive Multi-Object Search (HIMOS), a hierarchical, reinforcement learning-based approach to learn both exploration and manipulation skills and to reason at a high level about the required steps. Reinforcement learning has shown to work well in unexplored environments with high-dimensional observation spaces [5, 6]. We combine learned motions for local exploration in continuous action spaces [3] and frontier exploration for long-horizon exploration [4] together with mobile manipulation skills for object interactions [7]. We use semantic maps as the central memory component, which have shown to be an expressive and sample-efficient representation for these tasks [3] and design a high-level action space that exploits the acquired knowledge about the environment. By leveraging explored object instance locations as navigation waypoints, our approach efficiently learns these complex tasks from little data and consistently achieves success rates above 90% even as the number of target objects increases. By equipping all the low-level skills with mobility, we remove the “hand-off” problem in which subpolicies have to terminate in the initial set of the following skill [8, 9]. Lastly, we transfer the trained agent to the real world and demonstrate that it successfully accomplishes these tasks in a real office environment. In particular, we replace the subpolicies from simulation with unseen real-world variations and find that the policy is able to generalize to these unseen subpolicies and is robust to failures in their execution, making it highly modular and flexible for transfer. Finally, we present ablation studies to evaluate the impact of the main design decisions.
@@
We present a hierarchical, reinforcement learning-based approach that combines exploration and manipulation skills based on semantic knowledge and instance navigation points to efficiently solve these long-horizon tasks.
@@
We demonstrate the capabilities of this approach in both simulated and real-world experiments and show that, given accurate perception, its decision making achieves zero-shot transfer to the real world, unseen environments, unseen subpolicies, and is robust to unseen failures.
@@ II Related Work
Object search tasks: Exploration and the ability to find items of interest is a key requirement for a wide range of downstream tasks. Previous works have proposed methods to maximize coverage of explored space [10] and to find specific objects of interest based on vision [5], auditory signals [6, 11] or target object categories [12, 13]. In ordered multi-object search tasks, the agent has to find items in a fixed order in game environments [14] or realistic 3D apartments [1]. In unordered multi-object search, the agent simply has to find the target objects as fast as possible, irrespective of the order [2, 3]. We focus on this unordered task. As we aim to demonstrate our system on a real robot, we follow Schmalstieg et al. [3] and use the full continuous action space. This is in contrast to most previous works which focus on a simplified granular discrete action space. Existing search tasks assume that the desired goals can be freely reached by the agent. The interactive navigation task [15] relaxes this assumption by placing objects that the robot has to push away to reach the target. In contrast, we introduce an interactive search task for mobile robots equipped with a manipulator, that requires interaction with articulated objects to clear the path or reveal concealed objects. This requires integrating navigation and manipulation. Lastly, in contrast to most previous works, we demonstrate that our approach successfully transfers to the real world.
@@
Exploration requires both understanding and memorizing the seen environment and decision-making to explore the remaining space. Previous work has introduced both implicit and explicit memory mechanisms. Implicit memory agents either learn a direct end-to-end mapping from RGB-D images to actions or store embeddings of previous observations and retrieve them with an attention mechanism [2]. Other methods build explicit maps of the environment by projecting the RGB-D inputs into a global map. Commonly, this map is also annotated with semantic labels [14, 3, 1]. Further, combining short- and long-term exploration by learning an auxiliary prediction of the direction to the next closest object has proven to result in a strong performance in continuous action spaces [3]. We use this approach for low-level exploration. Frontier exploration [4] samples points on the frontier of the explored space and then navigates to these points. Instead of sampling, Ramakrishnan et al. [16] predict a potential function towards a target object. SGoLAM [17] combines mapping and a goal detection module. If no goal object is detected, it explores with frontier exploration and navigates directly to the goal otherwise. This results in strong results without any learning component. We include a similar exploration strategy as an option in our hierarchical approach. Zheng et al. [18, 19] leverage POMDP-solvers over an explicit belief representation to plan next viewpoints.
@@
Articulated object manipulation such as opening doors and drawers requires control of both the base and arm of the robot [20]. Existing approaches often separate both aspects and execute sequential navigation and manipulation. In our evaluation in simulation, we follow this approach and use BiRRT [21] to generate a motion plan for the robot arm. Recent works train an agent to control the base of the robot via reinforcement learning to follow given end-effector motions [22, 7]. We use this method in our real world evaluation as it generalizes across different robots, tasks, and environments.
@@
Hierarchical methods introduce layers of abstraction by decomposing the decision-making into higher and lower-level policies. This shortens the time horizon of the Markov decision process (MDP) for the higher-level policies and enables the agent to combine different modules or skills at lower levels. At the same time, joint training of low- and high-level policies is often unstable and hard to optimize [23]. We focus on combining pretrained subpolicies. Joint finetuning of these policies offers a path to further performance improvements in the future. Pretraining is a common approach to increase the stability of the policy in hierarchical reinforcement learning [24]. While naive skill-chaining of arbitrary skills often results in “hand-off” failures in which the subsequent skill cannot start from the current state [9], we resolve this issue by adding mobility to all the low-level skills in the real world execution, without the need for region-rewards [8]. A common navigation abstraction is to learn to set waypoints [25, 26], however, these often end up as very near points to the agent. In contrast, we propose instance navigation that provides a prior on important locations and action granularity. ASC [27] learns to combine navigation and pick skills for given receptacle locations. In contrast, we learn to search in unexplored environments with objects hidden in articulated objects. Alternatively, behavior trees are often used to decompose tasks hierarchically into a tree structure. These trees can either be fully constructed manually or be used in conjunction with a planner [28]. [29] learn values of tree nodes for cluttered object relocation. In contrast to these approaches, our proposed HIMOS learns high-level decision-making with a two-layer hierarchy.
@@ IV Hierarchical Interactive Multi-Object Search
The challenges introduced by this task require the robot to master an increasing number of different behaviors and skills. To stem this increase in complexity, we propose a hierarchical, reinforcement-based learning approach and design efficient abstractions over states and actions. The model consists of three main components: a mapping module, a set of subpolicies for exploration, navigation, and interaction, and a high-level policy module to select the next low-level action to execute. An overview of the approach is depicted in Figure 2.
Assumptions: The focus of this work is on learning long-horizon decision-making, exploration, and search. We abstract from real-world perception and assume to have access to the following: (i) a semantic camera that produces accurate semantic object labels, (ii) accurate depth and localization, (iii) an object detection module to perceive object poses in the environment and to detect whether an object interaction was successful. In training and simulation, the object detection module uses ground truth poses from the simulator while in real-world experiments, we rely on AR markers placed on the respective objects. Interaction failure is detected based on the change in pose of these markers after the interaction. (iv) the environment remains static except for the agent’s own interactions with it. For example, objects are not being displaced by other people during the search process.
@@ IV-A Mapping Module
A semantic map of the environment serves as a central memory component across all policies and modules. To construct this map, we build upon our previously introduced mapping module [3]. We extend this map to a partial panoptic segmentation [30] map by labeling task-specific objects such as doors and cabinets with an instance-specific color, which is randomly assigned whenever a new instance is detected. I.e. an instance’s color changes from episode to episode, but remains consistent within an episode. The robot receives the semantic masks and depth at a resolution of pixels, generates a point cloud from this information, and projects the points into a local top-down map, using the top-most point for each cell. With the local map and its current localization, the agent then updates an internal global map which is further annotated with the agent’s trace and encoded into an RGB image. Target objects are mapped as a special color, making the approach agnostic to the underlying classes of the target objects. After the agent has segmented and approached a target object, it updates the object’s annotation with a fixed color coding to mark the corresponding object as “found”. In the simulation, the robot receives the semantic labels from the simulator’s semantic camera. In the real world, we simulate access to semantic labels as follows: we pre-build a map of the environment using Hector SLAM [31] and annotate it with labels for target objects, cabinets, and doors. All the other occupied space is mapped to the wall category. At test time, we then use the robot’s depth camera to build the same local map as in the simulation and overlay it with this pre-annotated semantic map. Localization in the real world is done on the same pre-recorded map.
In each step, the agent then extracts an egocentric map from the global map and passes two representations of this map to the encoder: a coarse map of dimension at a resolution of and a fine-grained map of dimension at a resolution of . I.e., they cover and , respectively.
@@ IV-B Subpolicy Behaviors
The respective navigation goals are set to a pose slightly in front, or for the goal objects directly to the detected pose of the corresponding object. This navigation goal is then fed to an -planner which produces a trajectory at a resolution of . For training speed, we do not execute the full path in simulation, but set the robot’s base pose to the generated waypoints and only collect observations from these points. In the real world, we use the ROS navigation stack to move the robot to the goal. The policy may fail in some situations, for example, due to collision with obstacles or narrow doorways. In this case, the agent returns to the last feasible waypoint, and given the updated semantic map , the high-level policy has to make a new decision.
@@ IV-C High-level Decision Making
Efficient high-level decision making requires the right level of abstractions of states and actions. We hypothesize that object- and instance-level decision making is such an efficient level of abstraction for embodied search tasks. We design a high-level policy around this idea. In particular: (i) We propose an instance navigation subpolicy that leverages the agent’s accumulated knowledge about the environment. It provides a prior on important places and on the granularity of navigation points, making it data-efficient and well-optimizable. (ii) As objects are discrete instances, the resulting full action space remains discrete, avoiding the complexities of mixed action spaces. At the same time, all the subpolicies still act directly in continuous action spaces, allowing for direct transfer to real robotic systems. (iii) We abstract from reasoning about exact robot placements in the real world by shifting the responsibility of mobility into the subpolicies. This ensures that the subpolicies can start from a large set of initial positions, resolving the “hand-off” problem from naive skill-chaining [9] and strongly simplifies the learning process for the high-level policy. Furthermore, it enables us to change out the subpolicies to unseen subpolicies in the real world. (iv) We incorporate subpolicy failures into the training process, enabling the high-level policy to learn a retrial behavior if execution fails.
@@
Action Space: The high-level policy acts in a Semi-Markov Decision Process (SMDP) in which the actions model temporarily extended behaviors and act at irregular intervals [33]. It makes a decision whenever the last invoked subpolicy returns, successful or not. The high-level action space consists of (i) invoking the local exploration policy [3], (ii) invoking global frontier exploration [4], (iii) instance navigation with subsequent object interaction (if available), where instance IDs are selected through a one-hot vector mapping to fixed colors. The current task instantiation assumes a maximum of ten instances per episode, resulting in an overall 12-dimensional discrete action space.
@@
Adaptive Discounting: The high-level policy acts at irregular intervals, as the duration of the subpolicies varies largely. We correct for this time bias and accurately reflect the long-term consequences of actions with adaptive discounting.
@@
Invalid Action Masking: The availability of high-level actions varies with the state of the environment, e.g. navigating to an object instance depends on the instance being mapped, opening a cabinet is only possible if it is mapped and closed. We infer a valid actions vector from the agent’s observations and incorporate it into the training by masking out invalid actions as well as including it in the observation space of the agent. As a result, the agent can learn more effectively, speeding up the training process [32]. We implement this by replacing the logits of the invalid actions with a large negative number.
@@
Observation Space and Architecture: The observation space of the high-level policy consists of the coarse semantic map , the -dimensional target object vector and the -dimensional binary valid actions vector (see invalid action masking), where is the number of instances, and a 22-dimensional robot state vector . The robot state consists of linear and angular base velocities, the last low-level actions, sum of collisions over the last ten steps, a current collision flag, and a normalized history over the last 16 high-level actions taken. Following previous work [10, 3], the agent first encodes the coarse semantic map with a ResNet-18 [34]. Then it concatenates these features with the other, structured observations. The high-level policy is trained with Proximal Policy Optimization (PPO) [35]. We report hyperparametersand full network architectures in the supplementary material.